OpenStudy (anonymous):

What are the only possible solutions to the equation below? (x-6)^2=7 I'm thinking that it may be, X = /7 + 6 and X = -/7 +6 Or X = /7 -6 and X = /7 +6 It obviously can't be X = -/7 -6 PLEASE HELP!

HanAkoSolo (jamierox4ev3r):

\((x-6)^2\)+7=?

OpenStudy (anonymous):

No =7

OpenStudy (anonymous):

Crap, wrote it wrong, xD

HanAkoSolo (jamierox4ev3r):

oh i misunderstood okay thx for clarifying. Funny name btw....

OpenStudy (anonymous):

Thx. :3

HanAkoSolo (jamierox4ev3r):

\((x-6)^2=7\)

OpenStudy (anonymous):

Indeed...

HanAkoSolo (jamierox4ev3r):

\(x^2+36=7\)

HanAkoSolo (jamierox4ev3r):

x^2+36=7 -36 -36 x^2=-29

OpenStudy (anonymous):

It can't be X = -/7 -6 right? Please, look @ my question and see my guestimates. :)

OpenStudy (anonymous):

\[x=6\pm \sqrt{7}\]

OpenStudy (anonymous):

\[X=\sqrt{7} +6\] Is a correct one right?

OpenStudy (anonymous):

Guys, please help.

OpenStudy (johnweldon1993):

I would have just taken the square root at the beginning for this particular one... \(\large (x - 6)^2 = 7\) \(\large \sqrt{(x - 6)^2} = \sqrt{7}\) \(\large x - 6 = \pm\sqrt{7}\) \(\large x = 6\pm\sqrt{7}\) 2 solutions...

OpenStudy (anonymous):

Sooo, it's basically these solutions - \[X = \sqrt{7} -6 \] and \[X = \sqrt{7} +6\] That's how they are written...

OpenStudy (johnweldon1993):

No that is incorrect....remember when you take the square root of that 7...it is +/- √7....the way you have it written it is +/- 6 \(\large \sqrt{7} - 6 \ne 6 - \sqrt{7} \)

OpenStudy (johnweldon1993):

The correct way of writing them is \[\large x = 6 - \sqrt{7} \] and \[\large x = 6 + \sqrt{7} \]

OpenStudy (anonymous):

Then Apex writes them wrong. I'm trying my best here xD, that is how they are written in Apexvs.

OpenStudy (anonymous):

THose are solutions, correct?

OpenStudy (johnweldon1993):

Yes they are the correct solutions

OpenStudy (anonymous):

Then in other-words, the way that Apex writes them(my last reply) was correct, just in a different form?

OpenStudy (anonymous):

Yes that is basically it

OpenStudy (johnweldon1993):

I mean..mathematically ...incorrect.....but since that is how they write them...correct lol

OpenStudy (anonymous):

Alright, let's see... :)

OpenStudy (anonymous):

So, it's \[X = \sqrt{7} -6\] and \[X = \sqrt{7}+6\] not \[X = \sqrt{7}+6\] and \[X = -\sqrt{7} +6\]

OpenStudy (johnweldon1993):

THE SECOND CHOICES! lol that is how they would be written in the different form!! x = √7 + 6 x = -√7 + 6

OpenStudy (anonymous):

This is getting real confusing

HanAkoSolo (jamierox4ev3r):

ikr? :P

OpenStudy (anonymous):

And this is Algebra 1... (Supposedly) I'm in 9th grade moving on to 10th, we had a student teacher the whole damned year...

OpenStudy (johnweldon1993):

lol....okay What I wrote as the answer... \[\large 6 \pm \sqrt{7} \] this IS equal to \[\large \sqrt{7} + 6 \] and \[\large -\sqrt{7} + 6 \] lol ...all clear??

OpenStudy (anonymous):

So basically the second choices, correct?

OpenStudy (johnweldon1993):

Yes! lol

OpenStudy (anonymous):

Kk, sorry for being annoying... /Best-Response to John. :) CORRECT! /closed

OpenStudy (ybarrap):

@ObamaBinLlama says, "I'm thinking that it may be, X = /7 + 6 and X = -/7 +6" ... and you would be correct.

OpenStudy (anonymous):

Thx Barrap